OpenStudy (anonymous):

Help please :D ^ = to the power of. Simplify - 72a^5b^3 ---------------------------------- (divide) 84a^4b^2

OpenStudy (anonymous):

6ab/7

OpenStudy (anonymous):

72a^5b^3 / (84a^4b^2) One of the law of indices, a^m / a^n = a^(m - n). So, 72a^5b^3 / (84a^4b^2) = 6a^(5 - 4)b^(3 - 2) / 7 = 6ab / 7

OpenStudy (uri):